# Pareto-Conditioned Diffusion Models for Offline Multi-Objective Optimization

### v1

@@ 1 Introduction
The most dominant approach to offline MOO is model-based optimization (Kim et al., 2025), typically using a surrogate model (Trabucco et al., 2021; Yu et al., 2021; Yuan et al., 2023b; Qi et al., 2022; Chen et al., 2023; Daulton et al., 2020; 2021). The surrogate is trained on the static dataset to approximate the black-box objective and then used with multi-objective evolutionary algorithms (MOEAs) (Deb, 2001; Deb et al., 2002) to search for promising solutions. More recent approaches use generative models to directly synthesize solutions (Kim et al., 2023). For instance, ParetoFlow (Yuan et al., 2025) employs flow matching-based generative models guided by a multi-objective predictor to generate samples along the Pareto front. However, these methods still rely on surrogate predictors to guide the generation process, inheriting the same potential risks of exploiting inaccuracies in the surrogate model (Trabucco et al., 2021).
@@ 2 Related Work
Prior work in offline MOO has primarily focused on model-based approaches (Xue et al., 2024; Kim et al., 2025). These methods train a surrogate model on a static dataset to approximate the black-box objective function. Surrogates can take the form of Deep Neural Networks (DNNs) (Trabucco et al., 2021; Yu et al., 2021; Qi et al., 2022; Yuan et al., 2023b; Chen et al., 2023), which scale well to high-dimensional search spaces, or Gaussian Processes (GPs) (Daulton et al., 2020; 2021), which provide uncertainty estimates useful for exploration. Once the surrogate is trained, a separate search algorithm, most commonly MOEAs such as NSGA-II (Deb, 2001; Deb et al., 2002), is used to optimize the learned proxy to find promising solutions (Xue et al., 2024). The performance of this multi-stage pipeline, however, is fundamentally bottlenecked by the accuracy of the surrogate model (Trabucco et al., 2021).
@@
More recent approaches use generative models to directly synthesize solutions (Kim et al., 2023). The primary advantage of these methods is their ability to generate diverse designs without relying on a separate, explicit search algorithm (Kim et al., 2025). For instance, ParetoFlow (Yuan et al., 2025) and PGD-MOO (Annadani et al., 2025) employ generative models guided by a multi-objective predictor. Similarly, LaMBO (Stanton et al., 2022) integrates Bayesian optimization to steer its diffusion model, while MOGFNs (Jain et al., 2023) utilize generative flow networks often steered by scalarization functions. However, the effectiveness of these methods is still dependent on a surrogate predictor or a scalarization function to guide the generator (Yuan et al., 2025; Annadani et al., 2025).
@@ 3 Background
The goal of MOO is to approximate the PF with a diverse and well-distributed set of solutions. When evaluating the objective function is expensive, as in drug discovery or architecture search, offline MOO uses a pre-collected dataset , with , without direct access to during training (Xue et al., 2024; Kim et al., 2025). Offline methods rely solely on this static dataset , though a limited budget of new queries to may be allowed during evaluation. Solution quality is typically measured by the Hypervolume (HV) indicator, which is calculated relative to a reference point derived from the best-known solutions when the true PF is unknown (Zitzler & Thiele, 1998; Zitzler et al., 2007).
@@ 4 Method
Generating diverse set of condition points.
A crucial question in conditional sampling is how to select the target conditioning points, which we denote as . While simple tactics exist for single-objective optimization (Kumar & Levine, 2020), they are ineffective for MOO where the goal is to approximate the entire Pareto front. We therefore propose a two-stage procedure to generate a set of conditioning points , designed to be both high-quality and diverse. The procedure is described below briefly, with full pseudocode and explanation in Appendix A.9.
To collect the set of representative points, we propose to use a procedure motivated by the survival method in NSGA-III (Deb & Jain, 2014). Firstly, the objective space is partitioned by generating direction vectors via Riesz s-Energy method (Hardin & Saff, 2005). Secondly, the points in the offline dataset are ranked via non-dominated sorting, resulting in a list of fronts , where each front consists of set of points with same rank. Lastly, we iterate over the fronts to assign each point a direction vector . Initially, each direction vector gets assigned a point that is closest to it in perpendicular distance. Then, the rest of points in front are iterated over one-by-one, and assigned to a direction vector with least amount of points at the moment. This procedure is continued until (point, direction vector) pairs have been created.
After collecting the set of representative points, the points are extrapolated towards the assigned direction vector to improve their quality over the points in the offline dataset. However, in practice the number of solutions is often larger than the number of direction vectors . Therefore multiple points get extrapolated towards the same direction, which decreases the diversity of the condition points. To compact this, the final condition points are obtained by adding zero-mean Gaussian noise to the extrapolated points, akin to the Gaussian mutation operator used in evolutionary algorithms. This reweighted objective, paired with reference direction sampling forms the core of the proposed PCD.
@@ 5.1 Overview
Finally, we include generative model baselines. ParetoFlow (Yuan et al., 2025) is the only other generative model specifically developed for offline MOO that we are aware of and is our primary comparison222PGD-MOO (Annadani et al., 2025) is omitted because the preprint reports results only on a limited subset of synthetic tasks, and our attempts to reproduce the reported results using the authors’ code were unsuccessful.. While LaMBO (Stanton et al., 2022) is a related diffusion-based approach, it was used to generate some of the benchmark datasets themselves. Including LaMBO as a baseline could therefore introduce bias, so it is excluded. We also include a simple but strong baseline, (best), which represents the performance of the non-dominated set of solutions present in the original offline dataset.
@@ 5.3 Results
The results in Table 1 clearly demonstrate the effectiveness of our approach. PCD achieves the best overall rank, highlighting its consistent high performance across a diverse set of problems. This consistency is particularly noteworthy as these results were obtained using a single, untuned set of hyperparameters, underscoring the method’s robustness. Even in the MORL and MONAS task sets, where its performance is weaker relative to other categories, PCD still outperforms the other generative baseline, ParetoFlow (Yuan et al., 2025), by a considerable margin.
@@ 5.4 Ablation studies
We conduct an ablation study to validate the two primary components of PCD introduced in Section 4: the multi-objective reweighting scheme and the reference-direction mechanism for generating conditioning points. We compare three data processing techniques: 1) no processing (N/A), 2) simple pruning of low-quality points as in (Xue et al., 2024), and 3) our proposed data reweighting. Additionally, we compare three strategies for generating conditioning points: our proposed reference-direction (Ref. Dir.) mechanism, conditioning on the best set of points in the offline dataset (), and a simpler baseline (Ideal) that randomly samples non-dominated points from the dataset and interpolates them toward the ideal objective vector, a strategy common in offline single-objective optimization. Note that the full PCD method combines our proposed reweighting with the reference-direction mechanism.
@@ 8 Reproducibility Statement
Our implementation is available at https://github.com/jatan12/PCDiffusion.
@@ A.2 Reweighing sensitivity analysis
Based on our analysis in Section A.1, we hypothesized that the performance degradation on high-variance datasets like ZDT2 could be mitigated by adjusting the reweighting hyperparameters. Specifically, increasing the temperature parameter in Equation 6 should make the weighting distribution more uniform, effectively ”softening” the reweighting and preserving more of the learning signal from lower-quality samples. To validate this, we perform a sensitivity analysis on for the same three tasks: ZDT2, C10/MOP2, and MO-Hopper.
@@
This outcome confirms our intuition: the higher the variance in dataset quality, the higher temperature is required to achieve optimal performance. This suggests that for practical applications, could be tuned based on an initial analysis of the dataset’s quality distribution. We note that the other reweighting hyperparameter, , which controls smoothing over bin sizes, was observed to have a negligible impact on performance in preliminary experiments, and thus was not included in this analysis.
@@ A.3 Training details
We apply classifier-free guidance (CFG) dropout (Ho & Salimans, 2022) during training with probability 0.25 and use the same denoiser configuration for all tasks. We maintain an exponential moving average (EMA) of the model parameters (Karras et al., 2022) and use the EMA weights at sampling to improve stability (Karras et al., 2024). Sampling is performed with the stochastic SDE sampler of EDM using default hyperparameters, increasing the number of diffusion timesteps to 1024 to enhance sample quality. The default hyperparameters for all components of our PCD framework—including the denoiser network, the EDM sampler, our multi-objective reweighting scheme, and our conditioning point generation mechanism—are summarized in Table 3.
@@ A.4 Computational Cost
Recent work has established a strong theoretical connection between diffusion models (Ho et al., 2020; Song et al., 2021b) and flow matching (Lipman et al., 2023; Liu et al., 2023), showing that they can be unified under the framework of continuous-time generative models (Gao et al., 2024; Albergo et al., 2025). In particular, when using Gaussian probability flow paths, the sampling process for both frameworks is governed by an Ordinary Differential Equation (ODE) that maps a Gaussian noise distribution to the data distribution. Consequently, using a diffusion model with a deterministic ODE sampler is mathematically equivalent to the sampling process of standard Gaussian flow matching (Gao et al., 2024). The primary practical differences lie in the parameterization of the vector field, the noise schedule, and the weighting of the training objective (Gao et al., 2024). These design choices have been extensively studied within the diffusion context, most notably in the EDM paper (Karras et al., 2022). This work uses the parameterization derived in EDM.
@@ A.8 Case Study 2: Scalability to Many-Objective Problems
A.9 Detailed look at condition point generation
Section 4 introduced procedure generate set of diverse condition points by utilizing procedure similar to the survival selection from NSGA-III (Deb & Jain, 2014), which is displayed in Algorithm 2. We begin by generating direction vector via Riesz Energy method (Hardin & Saff, 2005), that partition the objective space . Then, our goal is to select a diverse subset of well-performing points from the objective space, and assign these points to one of the direction vectors. To do so, an empty set is initialized to hold the point, direction vector pairs. Then, a set of fronts is obtained via non-dominated sort. The procedure continues by iterating over each front . During each iteration, each direction is assigned a point with the shortest perpendicular distance to it (lines 8-13 in Alg. 2). Then, the rest of the points in (which have no direction vector assigned to them), are iterated over one-by-one, and assigned a direction vector with the least amount of points at the moment (lines 14-22 in Alg. 2).
After obtaining the set of (point, direction vector) pairs , the points are interpolated towards the corresponding direction vectors. Finally, Gaussian noise is added to the points to ensure further diversity.
The procedure described in Algorithm 2 follows very closely the survival procedure of NSGA-III with one key difference: in NSGA-III the points in the current population are assigned to the closest direction vector, and only if there is a need for more points, the points get assigned to the direction vectors with the least number of points. On the contrary, in our procedure, only points front get automatically assigned the direction vector closest to them. The rest of the points in the current front are assigned based on the number of points each direction vector has currently. This is crucial for ensuring that the conditioning points have an adequate coverage of the objective space in cases where the offline dataset itself has a poor coverage of the objective space.

### v2

@@ 1 Introduction
The most dominant approach to offline MOO is model-based optimization (Kim et al., 2026), typically using a surrogate model (Trabucco et al., 2021; Yu et al., 2021; Yuan et al., 2023b; Qi et al., 2022; Chen et al., 2023; Daulton et al., 2020; 2021). The surrogate is trained on the static dataset to approximate the black-box objective and then used with multi-objective evolutionary algorithms (MOEAs) (Deb, 2001; Deb et al., 2002) to search for promising solutions. More recent approaches use generative models to directly synthesize solutions (Kim et al., 2023). For instance, ParetoFlow (Yuan et al., 2025) employs flow matching-based generative models guided by a multi-objective predictor to generate samples along the Pareto front. However, these methods still rely on surrogate predictors to guide the generation process, inheriting the same potential risks of exploiting inaccuracies in the surrogate model (Trabucco et al., 2021).
@@ 2 Related Work
Prior work in offline MOO has primarily focused on model-based approaches (Xue et al., 2024; Kim et al., 2026). These methods train a surrogate model on a static dataset to approximate the black-box objective function. Surrogates can take the form of Deep Neural Networks (DNNs) (Trabucco et al., 2021; Yu et al., 2021; Qi et al., 2022; Yuan et al., 2023b; Chen et al., 2023), which scale well to high-dimensional search spaces, or Gaussian Processes (GPs) (Daulton et al., 2020; 2021), which provide uncertainty estimates useful for exploration. Once the surrogate is trained, a separate search algorithm, most commonly MOEAs such as NSGA-II (Deb, 2001; Deb et al., 2002), is used to optimize the learned proxy to find promising solutions (Xue et al., 2024). The performance of this multi-stage pipeline, however, is fundamentally bottlenecked by the accuracy of the surrogate model (Trabucco et al., 2021).
@@
More recent approaches use generative models to directly synthesize solutions (Kim et al., 2023). The primary advantage of these methods is their ability to generate diverse designs without relying on a separate, explicit search algorithm (Kim et al., 2026). For instance, ParetoFlow (Yuan et al., 2025) and PGD-MOO (Annadani et al., 2025) employ generative models guided by a multi-objective predictor. Similarly, LaMBO (Stanton et al., 2022) integrates Bayesian optimization to steer its diffusion model, while MOGFNs (Jain et al., 2023) utilize generative flow networks often steered by scalarization functions. However, the effectiveness of these methods is still dependent on a surrogate predictor or a scalarization function to guide the generator (Yuan et al., 2025; Annadani et al., 2025).
@@ 3 Background
The goal of MOO is to approximate the PF with a diverse and well-distributed set of solutions. When evaluating the objective function is expensive, as in drug discovery or architecture search, offline MOO uses a pre-collected dataset , with , without direct access to during training (Xue et al., 2024; Kim et al., 2026). Offline methods rely solely on this static dataset , though a limited budget of new queries to may be allowed during evaluation. Solution quality is typically measured by the Hypervolume (HV) indicator, which is calculated relative to a reference point derived from the best-known solutions when the true PF is unknown (Zitzler & Thiele, 1998; Zitzler et al., 2007).
@@ 4 Method
Generating a diverse set of conditioning points.
A crucial question in conditional sampling is how to select the target conditioning points, which we denote as . While simple tactics exist for single-objective optimization (Kumar & Levine, 2020), they are ineffective for MOO where the goal is to approximate the entire Pareto front. We therefore propose a two-stage procedure to generate a set of conditioning points , designed to be both high-quality and diverse. The procedure is described briefly below, with full pseudocode and explanation in Appendix A.9.
To collect the set of representative points, we propose to use a procedure motivated by the survival method in NSGA-III (Deb & Jain, 2014). Firstly, the objective space is partitioned by generating direction vectors via Riesz s-Energy method (Hardin & Saff, 2005). Secondly, the points in the offline dataset are ranked via non-dominated sorting, resulting in a list of fronts , where each front consists of a set of points with the same rank. Lastly, we iterate over the fronts to assign each point a direction vector . Initially, each direction vector gets assigned a point that is closest to it in perpendicular distance. Then, the rest of points in front are iterated over one-by-one and assigned to a direction vector with the least amount of points at the moment. This procedure is continued until (point, direction vector) pairs have been created.
After collecting the set of representative points, the points are extrapolated towards the assigned direction vector to improve their quality over the points in the offline dataset. However, in practice the number of solutions is often larger than the number of direction vectors . Therefore multiple points get extrapolated towards the same direction, which decreases the diversity of the conditioning points. To mitigate this, the final conditioning points are obtained by adding zero-mean Gaussian noise to the extrapolated points, akin to the Gaussian mutation operator used in evolutionary algorithms. This reweighted objective, paired with reference direction sampling, forms the core of the proposed PCD.
@@ 5.2 Baselines
Finally, we include generative model baselines. ParetoFlow (Yuan et al., 2025) is the only other generative model specifically developed for offline MOO that we are aware of and is our primary comparison222PGD-MOO (Annadani et al., 2025) is omitted because the results in the original study were limited to synthetic tasks, and we were unable to obtain results on our broader benchmarks using the official code.. While LaMBO (Stanton et al., 2022) is a related diffusion-based approach, it was used to generate some of the benchmark datasets themselves. Including LaMBO as a baseline could therefore introduce bias, so it is excluded. We also include a simple but strong baseline, (best), which represents the performance of the non-dominated set of solutions present in the original offline dataset.
@@ 5.3 Results
The results in Table 1 clearly demonstrate the effectiveness of our approach. PCD achieves the best overall rank, highlighting its consistent high performance across a diverse set of problems. This consistency is particularly noteworthy, as these results were obtained using a single, fixed set of hyperparameters, underscoring the method’s robustness in the offline setting where such tuning is inherently difficult. Even in the MORL and MONAS task sets, where its performance is weaker relative to other categories, PCD still outperforms the other generative baseline, ParetoFlow (Yuan et al., 2025), by a considerable margin.
@@ 5.4 Ablation studies
We conduct an ablation study to validate the two primary components of PCD introduced in Section 4: the multi-objective reweighting scheme and the reference-direction mechanism for generating conditioning points. We compare three data processing techniques: 1) no processing (N/A), 2) simple pruning of low-quality points as in (Xue et al., 2024), and 3) our proposed data reweighting. Additionally, we compare three strategies for generating conditioning points: our proposed reference-direction (Ref. Dir.) mechanism, conditioning on the best set of points in the offline dataset (), and a simpler baseline (Ideal) that randomly samples non-dominated points from the dataset and extrapolates them toward the ideal objective vector, a strategy common in offline single-objective optimization. Note that the full PCD method combines our proposed reweighting with the reference-direction mechanism.
@@ 8 Reproducibility Statement
Our implementation is available at https://github.com/jatan12/PCD.
@@ A.2 Reweighing sensitivity analysis
Based on our analysis in Section A.1, we hypothesized that the performance degradation on high-variance datasets like ZDT2 could be mitigated by adjusting the reweighting hyperparameters. Specifically, increasing the temperature parameter in Equation 6 should make the weighting distribution more uniform, effectively “softening” the reweighting and preserving more of the learning signal from lower-quality samples. To validate this, we perform a sensitivity analysis on for the same three tasks: ZDT2, C10/MOP2, and MO-Hopper.
@@
This outcome confirms our intuition: the higher the variance in dataset quality, the higher temperature is required to achieve optimal performance. Although we explored systematic ways to tune (e.g., estimating it from the standard deviation or CDF of the dominance number distribution), we found the default value () to be robust across the majority of tasks. However, for outlier datasets with high variance like ZDT2, we recommend a simple visual heuristic: if the distribution of dominance numbers exhibits a “long tail,” increasing effectively “softens” the reweighting to preserve more training data. We note that the other reweighting hyperparameter, , which controls smoothing over bin sizes, was observed to have a negligible impact on performance in preliminary experiments, and thus was not included in this analysis.
@@ A.3 Training details
We apply CFG dropout (Ho & Salimans, 2022) during training with probability 0.25 and use the same denoiser configuration for all tasks. We maintain an exponential moving average (EMA) of the model parameters (Karras et al., 2022) and use the EMA weights for sampling to improve stability (Karras et al., 2024). Sampling is performed with the EDM stochastic differential equation (SDE) sampler using default hyperparameters, increasing the number of diffusion timesteps to 1024 to enhance sample quality. The default hyperparameters for all components of our PCD framework—including the denoiser network, the EDM sampler, our multi-objective reweighting scheme, and our conditioning point generation mechanism—are summarized in Table 3.
@@ A.4 Computational Cost
Recent work has established a strong theoretical connection between diffusion models (Ho et al., 2020; Song et al., 2021b) and flow matching (Lipman et al., 2023; Liu et al., 2023), showing that they can be unified under the framework of continuous-time generative models (Gao et al., 2025; Albergo et al., 2025). In particular, when using Gaussian probability flow paths, the sampling process for both frameworks is governed by an ODE that maps a Gaussian noise distribution to the data distribution. Consequently, using a diffusion model with a deterministic ODE sampler is mathematically equivalent to the sampling process of standard Gaussian flow matching (Gao et al., 2025). The primary practical differences lie in the parameterization of the vector field, the noise schedule, and the weighting of the training objective (Gao et al., 2025). These design choices have been extensively studied within the diffusion context, most notably in the EDM paper (Karras et al., 2022). This work uses the parameterization derived in EDM.
@@ A.8 Case Study 2: Scalability to Many-Objective Problems
A.9 Detailed look at conditioning point generation
Section 4 introduced the procedure to generate a set of diverse conditioning points by utilizing a procedure similar to the survival selection from NSGA-III (Deb & Jain, 2014), which is detailed in Algorithm 2. We begin by generating direction vectors via Riesz s-Energy method (Hardin & Saff, 2005), which partition the objective space . Next, our goal is to select a diverse subset of well-performing points from the objective space and assign these points to one of the direction vectors. To do so, an empty set is initialized to hold pairs of points and direction vectors . Then, a set of fronts is obtained via non-dominated sorting. The procedure continues by iterating over each front . During each iteration, we identify the point with the shortest perpendicular distance to each direction vector and assign it to that vector (lines 8-13 in Alg. 2). Finally, the remaining points in (which have no direction vector assigned to them), are iterated over one-by-one and assigned to the direction vector that currently has the fewest assigned points (lines 14-22 in Alg. 2).
After obtaining the set of (point, direction vector) pairs , the points are extrapolated towards the corresponding direction vectors. Finally, Gaussian noise is added to the points to ensure further diversity. The procedure described in Algorithm 2 follows very closely to the survival procedure of NSGA-III with one key difference: in NSGA-III the points in the current population are assigned to the closest direction vector, and only if there is a need for more points, the points get assigned to the direction vectors with the least number of points. On the contrary, in our procedure, only points from a front get automatically assigned the direction vector closest to them. The rest of the points in the current front are assigned based on the number of points each direction vector has currently. This is crucial for ensuring that the conditioning points have an adequate coverage of the objective space in cases where the offline dataset itself has a poor coverage of the objective space.